If the speed of the bob at the lowest point A is 0.84 m/s, what is its KE at this point?


1
Expert's answer
2022-04-21T13:01:28-0400

"KE=\frac{mv^2}2=\frac{m\cdot 0.84^2}2=0.35m ~(\text J)."